OpenStudy (rosales0228):

WILL GIVE MEDAL!! You invest 3,200 in an account that pays interest rate of %7.25, compound contentiously. Calculate the balance of you account after 12 yrs. Round your answer to the nearest hundredth. I got 272471, is that correct??

OpenStudy (rosales0228):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

A = P*e^(r*t) A = 3200*e^(0.0725*12) A = 7,638.11473128001 A = 7,638.11 So the balance in 12 yrs is $7,638.11

jimthompson5910 (jim_thompson5910):

not sure how you got that number, but it's definitely way too high

OpenStudy (rosales0228):

i dont understand what i am doing wrong:/

jimthompson5910 (jim_thompson5910):

how did you type it in?

OpenStudy (rosales0228):

3200*e^7.025*12

jimthompson5910 (jim_thompson5910):

i see, that 7.025 should be 0.0725

OpenStudy (help!!!!):

AI=P+PRT

jimthompson5910 (jim_thompson5910):

and make sure you surround 0.0725*12 in parenthesis

OpenStudy (rosales0228):

okay thank you!

jimthompson5910 (jim_thompson5910):

np